OpenStudy (anonymous):

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the line? Answer -3/5 1/3 3 -3

OpenStudy (anonymous):

http://www.twiddla.com/1027275

OpenStudy (precal):

find you slope first then take the opposite sign and the reciprocal of the number

OpenStudy (anonymous):

What is the anwser?

OpenStudy (precal):

so what is the slope?

OpenStudy (precal):

gotta go, maybe someone will help you

OpenStudy (anikay):

Ok, so you need to solve for y, then do as precal said and take the opposite reciprocal meaning negative flipped fraction

OpenStudy (anikay):

so you got?

OpenStudy (anonymous):

nope!

OpenStudy (precal):

take your equation and solve for y tell me what equation you get at that point hint: y=?

OpenStudy (anonymous):

No idea! I am So stupid!

OpenStudy (anikay):

Sorry, once we get down to pre-algebra stuff I forget how to teach it. :/ Best I can say is whatever you do to one side you have to do it to the other side.

OpenStudy (precal):



OpenStudy (anonymous):

Write the slope-intercept equation of the line perpendicular to y = 5/2 x - 2, which passes though the point (0, 2). Answer y = 5x + 2 y = -2/5x + 2 y = 2x + 5 y = 1/2x + 5

OpenStudy (precal):

the number in front of x is (-1/3) you want +3

OpenStudy (precal):

sorry gotta go

OpenStudy (anonymous):

First change it into Slope-Intercept form. y = mx + b where m is slope. x + 3y = 5 Subtract x. 3y = -x + 5 Divide by 3. y = -1/3 + 5/3 So your slope is -1/3. However it asks for the slope of the line perpendicular to this one. Slope of perpendicular lines are the opposite reciprocal - change the value from positive to negative or vice versa, then find the reciprocal. ex. opposite reciprocal of -1/2 is +2/1 or 2

OpenStudy (anonymous):

So what is the opposite reciprocal of -3/1?

OpenStudy (anonymous):

1/3!

OpenStudy (anonymous):

Sorry I meant to say, what is the opposite reciprocal of -1/3. My mistake!

OpenStudy (anonymous):

So what is the opposite reciprocal of -1/3?

OpenStudy (precal):

I believe I stated it already +3

OpenStudy (anonymous):

ohh sorry! Didn't see it. (: